OpenStudy (anonymous):

a walks at uniform rate of 4km an hour 4 hours after his start b cycles after him at 10kmph how far wil b catch a after d starting point

OpenStudy (anonymous):

sum1 help

OpenStudy (anonymous):

do u knw it

OpenStudy (anonymous):

look at the units they might help. A travels at 4 [km/hour] and A travels for 4 hours So, aldren, how far does A travel total in that amount of time?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

alright... great job there. So A traveled for 16 km. Now, how long does it take for B to travel the same amount of distance?

OpenStudy (anonymous):

160 mins

OpenStudy (anonymous):

hhmmm you wouldn't get minutes. remember our problem is in hours so far

OpenStudy (anonymous):

in 4 hrs a has gone a distance of 4*4=16km now since they r moving in same direction their relative speed =10-4=6 km/hr hence required time is 16/6 =8/3 =2 hr 40 mins hence distance from starting point =16 + 10*8/3=16+80/3 =42.66 (appx) from starting point..

OpenStudy (anonymous):

2hrs 40mins

OpenStudy (anonymous):

kwo8 can u explain

OpenStudy (anonymous):

matricked ur answer is wrong

OpenStudy (anonymous):

i dint undrstand

OpenStudy (ajprincess):

velocity=distance/time distance=time*velocity The distance covered by a in t hours=The distance covered by b in t-4 hours t*4=(t-4)*10 4t=10t-40 40=10t-4t 40=6t 40/6=t d=40/6*4 =160/6 =26.67km is it clear?

OpenStudy (anonymous):

@aldren may i know the ans ..

OpenStudy (ajprincess):

ur answer is perfect @matricked. Since the question is hw far we need to find the distance.